# Systematic biases from the exclusion of higher harmonics in parameter estimation on LISA binaries

### v1

@@ I Introduction
As the catalog of gravitational wave (GW) events grows and the detectors improve in sensitivity, our ability to characterize the properties of individual events and of compact binary populations will be increasingly limited by systematic effects. Some of these systematic effects are due to detector noise and astrophysics, but waveform systematics are particularly important, and they have attracted a significant amount of attention in the context of ground-based detectors (see e.g. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]).
@@
The Laser Interferometer Space Antenna (LISA) is expected to observe binary black holes (BBHs) with much larger total masses and mass ratios, and with considerably louder signal-to-noise ratios (SNRs), than anything we have seen thus far in ground-based observatories (see e.g. Refs. [16, 17]). It has long been recognized that systematic biases in waveform modeling could be considerably worse for parameter estimation (PE) with signals detected by LISA [18, 19, 20, 21]. Thus far, however, little work has been done to attempt to understand just how significant these biases will be, and how they may vary across the parameter space.
@@
One source of systematic bias expected to affect PE with LISA sources is the bias due to neglecting higher harmonics in waveforms. It has been known for some time that more energy is radiated in higher multipoles for systems with larger mass ratios [22], and that the inclusion of higher multipoles will significantly enhance the science return of ringdown observations with LISA [23, 24, 25].
@@
Numerous authors have studied the biases introduced by neglecting higher-order modes in PE on BBH signals in current ground-based observatories, generally finding significant systematic errors due to the omission of higher modes for systems with high mass ratio, large total mass, high SNR, and inclination angle close to $\pi/2$ (edge-on systems) [26, 27, 28, 29, 30, 31, 32]. Some of these and other works have additionally found that relying on quadrupole-only waveforms decreases the efficacy of searches for BBH systems in current ground-based detectors, again in particular for heavier, more asymmetric, and more edge-on binaries [33, 34, 35]. A recent series of papers found that a combination of downweighting glitches and including higher modes in searches of the data of the third observing run of LVK resulted in 14 new BBH event candidates, in addition to increasing the significance of several candidates previously deemed “marginal” [36, 37, 38], as well as increasing the overall sensitive volume of GW searches [39]. On the other hand, another recent study found that the exclusion of higher harmonics caused relatively minimal biases in parametrized post-Einsteinian (ppE) tests of general relativity, at least compared to the biases induced by other neglected physics (i.e., precession) [15]. However, it is unclear how this result would extend to considerably more massive events. There is clearly motivation to account for higher modes more carefully, both in detection and in PE with GW signals. In this study, we will focus on the importance of subdominant modes in performing PE on BBHs observed by LISA.
@@
One previous study [40] demonstrated that, when performing PE on a high-SNR massive black hole binary (MBHB) event of the kind anticipated to be observed with LISA, systematic biases arise due to insufficient mode content in the waveform templates used for signal recovery. This phenomenon was demonstrated explicitly in Ref. [40] for a single example binary. In this paper, we perform the same analysis of varying the mode content in waveform templates for several example MBHB systems, studying how the aforementioned systematic biases in PE change for events with different total masses, mass ratios, and inclination angles. We then investigate the extent to which these biases can be predicted in a cost-efficient manner, i.e., without having to perform PE for every MBHB event under consideration. We use one such method (direct likelihood optimization; see Sec. II.3), which we find to be remarkably accurate in estimating systematic biases, to set approximate boundaries on the parameter space in which unbiased PE can be performed (i.e., where there is no significant bias due to neglecting a higher-order mode in the waveform template).
@@
where $m_{1},m_{2}$ and $\chi_{1},\chi_{2}$ are the individual masses and dimensionless spins of the progenitor BHs, respectively. For each bin in the grid of detector-frame total mass $M/M_{\odot}\in[3\times 10^{5},10^{6}]$ and mass ratio $q\in[1.1,10]$, we show the minimum redshift at which the systematic bias on parameters [$\mathcal{M}_{c},q,\chi_{+},\chi_{-}$] due to excluding the $(\ell,~{}m)=(3,~{}2)$ mode is still less than the $2\sigma$ statistical error on the parameter. Results are shown for systems with inclination angle $\iota=\pi/3$. Moving from left to right in the plot, we see that biases initially become worse as we move toward more massive and more asymmetric binaries, as expected. Continuing further, we have the balancing effect of the chirp mass decreasing going from top left to bottom right, such that mergers occur at higher frequencies. When the merger, where higher-order modes are most important, is pushed toward the higher-frequency region where LISA is less sensitive, the bias due to neglecting these modes becomes less severe again. The reason for the chosen range of total mass ($y$-axis) will be clear in Sec. III, where we find that our PE results on MBHB events with total mass $\lesssim 3\times 10^{5}\,M_{\odot}$ do not exhibit significant biases, whereas the PE for events with total mass $\sim 10^{6}\,M_{\odot}$ is significantly biased. Details on the construction of Fig. 1, as well as similar plots for extrinsic parameters and for systems with inclination $\iota\approx\pi/2$, are given in Sec. V.2.
@@
Figure 1: Minimum redshift at which parameter estimation is unbiased, in the sense that the systematic bias on all four intrinsic parameters $\mathcal{M}_{c},\,q,\,\chi_{+},$ and $\chi_{-}$ due to neglecting the $(\ell,m)=(3,~{}2)$ mode in the waveform template is less than the $2\sigma$ statistical error on the parameters. For example, for $M=7\times 10^{5}\,M_{\odot}$ and $q=2$ the parameter estimation is biased when $z\lesssim 3$. The grid is log-spaced between $q\in[1.1,3]$ and spaced linearly between $q\in[3,10]$ (with the transition marked by the dashed gray line), due to the more significant changes in results observed as we move from nearly symmetric to clearly asymmetric binaries. The corresponding plot for extrinsic parameters (sky localization and distance) is given in Fig. 14.
@@
In the following sections, we lay out the components that go into the construction and interpretation of Fig. 1. In Sec. II, we outline how we perform both full PE with lisabeta [41] and several cheaper methods of estimating systematic bias. In Secs. III and IV, we show the results of PE for the intrinsic and extrinsic parameters of about 20 selected binaries, respectively. In Sec. V, we show how our methods of rapidly approximating bias compare to PE and present more plots similar to Fig. 1. Finally, in Sec. VI, we show the results of performing PE with crude waveforms containing more higher-order modes than are currently modeled with the waveform approximants we use for PE (IMRPhenomXHM [42]). Throughout this paper we use geometrical units ($G=c=1$).
@@ II.1 Parameter estimation with lisabeta
We examine the extent of systematic biases due to incomplete mode content in the GW templates used for PE on MBHBs of the kind expected to be observed by LISA. We use lisabeta [41] to perform PE varying the redshifted total mass ($M/M_{\odot}=10^{5},10^{6}$), mass ratio ($q=1.1,4,8$), and inclination angle ($\iota=\pi/12,\pi/3,\pi/2-\pi/12$). Observing stark differences in the results for events with total mass $10^{5}\,M_{\odot}$ vs. $10^{6}\,M_{\odot}$, we also run PE on a few additional events at the intermediate mass of $3\times 10^{5}\,M_{\odot}$, with the median inclination angle ($\iota=\pi/3$). Altogether, this amounts to a total of 21 PE runs with lisabeta. The rest of the parameters (not modified between runs) are listed in Table 1, where $\chi_{1,2}$ are the dimensionless progenitor spins aligned with the orbital angular momentum of the binary; $D_{L}$ is the luminosity distance in Mpc, which we choose to correspond to redshift $z=1$; and $t_{c}$ is the time of coalescence. The angles $\beta_{L}$ and $\lambda_{L}$ are the ecliptic latitude and longitude, respectively, in the LISA frame; $\phi$ is BBH source frame phase; and $\Psi_{L}$ is the polarization in the LISA frame.
@@
For each combination of parameters, we generate the injected waveform for the event with the IMRPhenomXHM waveform model [42, 43], including the $(\ell,~{}m)=$ (2, 2), (2, 1), (3, 3), (3, 2) and (4, 4) modes – i.e., all the available modes with the IMRPhenomXHM waveform approximant. (From here on, whenever we denote a mode by a comma-delimited pair of numbers in parentheses, we will be referring to the angular harmonic indices of the mode, as done here.) We then perform PE on the signal with lisabeta using first a template with all five modes, then we iterate with one less mode, until we are attempting to recover the signal with just the quadrupole. The authors of Ref. [40] chose to “deactivate” modes in order of SNR contribution, such that a four-mode template excludes only the mode with the lowest SNR, the three-mode template excludes the two “quietest” modes, etc. To simplify comparisons between our different events, we always include modes in the following order:
@@
This ordering was chosen for a number of reasons.
First, when examining the mode-by-mode SNR of the 21 events we considered, 9 out of the 21 exhibited the above ordering. In Fig. 2, we can see the prevalence of this mode ordering within the region of parameter space we consider. For each inclination angle, different colors indicate a different ordering of the subdominant modes (ranked from “loudest” to “quietest”). In all the configurations shown here, the (2, 2) mode dominates, with the (3, 3) mode generally being the second loudest. The ordering of the (2, 1), (4, 4), and (3, 2) modes varies a bit more, with the (4, 4) usually being the quietest for a nearly face-on system, and the (3, 2) generally being the quietest otherwise.
@@
We use the power spectral density (PSD) required by the LISA Science Requirements Document (SciRDv1) [44]. We assume the frequency range of LISA to be $10^{-5}-0.5\,{\rm Hz}$ and incorporate the full response of LISA [45]. We also account for a background of white dwarf noise as in Sec. 9 of Ref. [44].
To sample the GW likelihood, we use the parallel-tempered Markov chain Monte Carlo (MCMC) ensemble sampler ptemcee [46, 47, 48], which is a fast sampler well-adapted to handling complex posteriors. To improve the robustness for degenerate posteriors, the ensemble sampler is enriched with jump proposals designed to explore the possible sky multimodality [41]. For a typical PE run, we use 64 walkers and run ptemcee at 10 temperatures with a max temperature of 100. For each event, we sample in the combinations of mass and spin parameters [$\mathcal{M}_{c},\,q,\,\chi_{+},$ and $\chi_{-}$] defined in Eq. (I). At first order in post-Newtonian (PN) theory, the GW frequency evolution is solely dependent on the chirp mass $\mathcal{M}_{c}$, so this is often the mass parameter that is most easily determined from the data. Similarly, $\chi_{+}$ (also commonly known as $\chi_{\rm{eff}}$) is generally much easier to determine from the data than the individual spins $\chi_{1}$ and $\chi_{2}$. Lastly, $q$ and $\chi_{-}$ are chosen to complete the set of mass and spin parameters, respectively.
To perform PE, we first calculate the Fisher covariance matrix and initialize chains from the results of this calculation. We set uniform priors on the chirp mass and distance in the range of 0.1 to 10 times their injected values. For the mass ratio, we use a uniform prior over the range [1, 10]110[1,\,10][ 1 , 10 ], and the spin parameters have a uniform prior over their allowed ranges of [1, 1]11[-1,\,1][ - 1 , 1 ].
@@ II.2 Predicting biases in the linear signal approximation
Given the computational cost of full Bayesian PE analyses, it is desirable to have a much faster, cheaper method of predicting systematic biases due to, e.g., neglecting higher-order modes. One technique for accomplishing this is to estimate systematic errors in the linear signal approximation as done by Cutler & Vallisneri (henceforth, “CV”) in Ref. [19]. In Ref. [49], the errors calculated in this approximation were found to reliably predict the biases in parameter inference of ET and LISA sources due to overlapping signals, foregrounds created by unfitted sources, and incorrectly removed sources, among other things. The authors of Ref. [40] examined the usefulness of this technique for the example binary they considered, finding that the errors were not properly reconstructed within the linear signal approximation. Here we expand the study of the usefulness of the CV estimates, covering a much larger region of parameter space than previously covered in Ref. [40]. in App. B, for completeness, we review the CV formalism. In Sec. V below we compare our PE results with the errors predicted by the CV estimates.

### v2

# Systematic biases in parameter estimation on LISA binaries: The effect of excluding higher harmonics for non-spinning binaries

@@ I Introduction
As the catalog of gravitational wave (GW) events grows and the detectors improve in sensitivity, our ability to characterize the properties of individual events and of compact binary populations will be increasingly limited by systematic effects. Some of these systematic effects are due to detector noise and astrophysics, but waveform systematics are particularly important, and they have attracted a significant amount of attention in the context of ground-based detectors (see e.g. [Purrer:2019jcp, Jan:2020bdz, Moore:2021eok, Hu:2022rjq, Hu:2022bji, Read:2023hkv, Owen:2023mid, Puecher:2023rxw, Jan:2023raq, Kapil:2024zdn, Dhani:2024jja, Gupta:2024gun, Pompili:2024yec, Bachhar:2024olc, Chandramouli:2024vhw, Khan:2024whs]).
@@
The Laser Interferometer Space Antenna (LISA) is expected to observe binary black holes (BBHs) with much larger total masses and mass ratios, and with considerably louder signal-to-noise ratios (SNRs), than anything we have seen thus far in ground-based observatories (see e.g. Refs. [LISA:2017pwj, LISA:2024hlh]). It has long been recognized that systematic biases in waveform modeling will consequently be considerably worse for parameter estimation (PE) with signals detected by LISA [Berti:2006ew, Cutler:2007mi, Ferguson:2020xnm, LISAConsortiumWaveformWorkingGroup:2023arg]. Although early estimates were already able to demonstrate that systematic errors would dominate over statistical errors for massive LISA sources [Littenberg:2012uj], thus far, relatively little detailed work has been done to attempt to understand just how significant these biases will be, and how they may vary across the parameter space.
@@
One source of systematic bias expected to affect PE with LISA sources is the bias due to neglecting higher harmonics in waveforms. It has been known for some time that more energy is radiated in higher multipoles for systems with larger mass ratios [Berti:2007fi], and that the inclusion of higher multipoles will significantly enhance the science return of ringdown observations with LISA [Baibhav:2017jhs, Baibhav:2018rfk, Baibhav:2020tma].
@@
Numerous authors have studied the biases introduced by neglecting higher-order modes in PE on BBH signals in current ground-based observatories, generally finding significant systematic errors due to the omission of higher modes for systems with high mass ratio, large total mass, high SNR, and inclination angle close to $\pi/2$ (edge-on systems) [Varma:2016dnf, CalderonBustillo:2015lrt, Littenberg:2012uj, Varma:2014jxa, Graff:2015bba, Shaik:2019dym, Pang:2018hjb]. Some of these and other works have additionally found that relying on quadrupole-only waveforms decreases the efficacy of searches for BBH systems in current ground-based detectors, again in particular for heavier, more asymmetric, and more edge-on binaries [Harry:2017weg, CalderonBustillo:2016rlt, Brown:2012nn]. A recent series of papers found that a combination of downweighting glitches and including higher modes in searches of the data of the third observing run of LVK resulted in 12 new BBH event candidates, in addition to increasing the significance of several candidates previously deemed “marginal” [Wadekar:2023gea, Wadekar:2023kym, Wadekar:2024zdq], as well as increasing the overall sensitive volume of GW searches [Mehta:2025jiq]. On the other hand, another recent study found that the exclusion of higher harmonics caused relatively minimal biases in parametrized post-Einsteinian (ppE) tests of general relativity, at least compared to the biases induced by other neglected physics (i.e., precession) [Chandramouli:2024vhw]. However, it is unclear how this result would extend to considerably more massive events. There is clearly motivation to account for higher modes more carefully, both in detection and in PE with GW signals. In this study, we will focus on the importance of subdominant modes in performing PE on BBHs observed by LISA.
@@
One previous study [Pitte:2023ltw] demonstrated that, when performing PE on a high-SNR massive black hole binary (MBHB) event of the kind anticipated to be observed with LISA, systematic biases arise due to insufficient mode content in the waveform templates used for signal recovery. This phenomenon was demonstrated explicitly in Ref. [Pitte:2023ltw] for a single example binary. In this paper, we perform the same analysis of varying the mode content in waveform templates for several example MBHB systems, studying how the aforementioned systematic biases in PE change for events with different total masses, mass ratios, and inclination angles. We then investigate the extent to which these biases can be predicted in a cost-efficient manner, i.e., without having to perform PE for every MBHB event under consideration. We use one such method (direct likelihood optimization; see Sec. II.3), which we find to be remarkably accurate in estimating systematic biases, to set approximate boundaries on the parameter space in which unbiased PE can be performed (i.e., where there is no significant bias due to neglecting a higher-order mode in the waveform template).
@@
where $m_{1},m_{2}$ and $\chi_{1},\chi_{2}$ are the individual masses and dimensionless spins of the progenitor BHs, respectively. For each bin in the grid of detector-frame total mass $M/M_{\odot}\in[3\times 10^{5},10^{6}]$ and mass ratio $q\in[1.1,10]$, we show the minimum redshift at which the systematic bias on parameters $\mathcal{M}_{c},q,\chi_{+},\chi_{-}$] due to excluding the $(\ell,~m)=(3,~2)$ mode is still less than the $2\sigma$ statistical error on the parameter. Results are shown for systems with inclination angle $\iota=\pi/3$. Moving from left to right in the plot, we see that biases initially become worse as we move toward more massive and more asymmetric binaries, as expected. Continuing further, we have the balancing effect of the chirp mass decreasing going from top left to bottom right, such that mergers occur at higher frequencies. When the merger, where higher-order modes are most important, is pushed toward the higher-frequency region where LISA is less sensitive, the bias due to neglecting these modes becomes less severe again. The reason for the chosen range of total mass ($y$-axis) will be clear in Sec. III, where we find that our PE results on MBHB events with total mass $\lesssim 3\times 10^{5}\,M_{\odot}$ do not exhibit significant biases, whereas the PE for events with total mass $\sim 10^{6}\,M_{\odot}$ is significantly biased. Details on the construction of Fig. 1, as well as similar plots for extrinsic parameters and for systems with inclination $\iota\approx\pi/2$, are given in Sec. V.2.
@@
Figure 1: Minimum redshift at which parameter estimation is unbiased, in the sense that the systematic bias on all four intrinsic parameters $\mathcal{M}_{c},\,q,\,\chi_{+},$ and $\chi_{-}$ due to neglecting the $(\ell,|m|)=(3,~2)$ mode in the waveform template is less than the $2\sigma$ statistical error on the parameters. For example, for $M=7\times 10^{5}\,M_{\odot}$ and $q=2$, the parameter estimation is biased when $z\lesssim 3$. The grid is log-spaced between $q\in[1.1,3]$ and spaced linearly between $q\in[3,10]$ (with the transition marked by the dashed gray line), due to the more significant changes in results observed as we move from nearly symmetric to clearly asymmetric binaries. The corresponding plot for extrinsic parameters (sky localization and distance) is given in Fig. 14.
@@
In the following sections, we lay out the components that go into the construction and interpretation of Fig. 1. In Sec. II, we outline how we perform both full PE with lisabeta [Marsat:2020rtl] and several cheaper methods of estimating systematic bias. In Secs. III and IV, we show the results of PE for the intrinsic and extrinsic parameters of about 20 selected binaries, respectively. In Sec. V, we show how our methods of rapidly approximating bias compare to PE and present more plots similar to Fig. 1. Finally, in Sec. VI, we show the results of performing PE with crude waveforms containing more higher-order modes than are currently modeled with the waveform approximants we use for PE (IMRPhenomXHM [Garcia-Quiros:2020qpx]). Throughout this paper we use geometrical units ($G=c=1$).
@@ II.1 Parameter estimation with lisabeta
We examine the extent of systematic biases due to incomplete mode content in the GW templates used for PE on MBHBs of the kind expected to be observed by LISA. We use lisabeta [Marsat:2020rtl] to perform PE varying the redshifted total mass ($M/M_{\odot}=10^{5},10^{6}$), mass ratio ($q=1.1,4,8$), and inclination angle ($\iota=\pi/12,\pi/3,\pi/2-\pi/12$). Observing stark differences in the results for events with total mass $10^{5}\,M_{\odot}$ vs. $10^{6}\,M_{\odot}$, we also run PE on a few additional events at the intermediate mass of $3\times 10^{5}\,M_{\odot}$, with the median inclination angle ($\iota=\pi/3$). Altogether, this amounts to a total of 21 PE runs with lisabeta. The rest of the parameters (not modified between runs) are listed in Table 1, where $\chi_{1,2}$ are the dimensionless progenitor spins aligned with the orbital angular momentum of the binary; $D_{L}$ is the luminosity distance in Mpc, which we choose to correspond to redshift $z=1$; and $\delta t$ is defined through $t_{c}=t_{0}+\delta t$, with $t_{c}$ the coalescence time and $t_{0}$ the reference epoch (in years). The angles $\beta_{L}$ and $\lambda_{L}$ are the ecliptic latitude and longitude, respectively, in the LISA frame; $\phi$ is the BBH source frame phase; and $\Psi_{L}$ is the polarization in the LISA frame. Note that while we vary the total mass between runs, we set our sampler to infer the chirp mass, $\mathcal{M}_{c}$, along with the mass ratio, $q$, both of which are defined in Eq. (I). Similarly, we sample in spin parameters $\chi_{+}$ and $\chi_{-}$, rather than $\chi_{1}$ and $\chi_{2}$, although we note that the injected values of these are all zero in the scope of this paper. Altogether, then, the 11 parameters that are simultaneously sampled in each PE run are $\mathcal{M}_{c},q,\chi_{+},\chi_{-},D_{L},\delta t,\beta_{L},\lambda_{L},\phi,\iota,\Psi_{L}$].
@@
For each combination of parameters, we generate the injected waveform for the event with the IMRPhenomXHM waveform model [Garcia-Quiros:2020qpx, Pratten:2020fqn], including the $(\ell,~m)=$ (2, 2), (2, 1), (3, 3), (3, 2) and (4, 4) modes – i.e., all the available modes with the IMRPhenomXHM waveform approximant. (From here on, whenever we denote a mode by a comma-delimited pair of numbers in parentheses, we will be referring to the angular harmonic indices of the mode, as done here.) We then perform PE on the signal with lisabeta using first a template with all five modes, then we iterate with one less mode, until we are attempting to recover the signal with just the quadrupole. The authors of Ref. [Pitte:2023ltw] chose to “deactivate” modes in order of SNR contribution, such that a four-mode template excludes only the mode with the lowest SNR, the three-mode template excludes the two “quietest” modes, etc. To simplify comparisons between our different events, we always include modes in the following order:
@@
This ordering was chosen for a number of reasons. First, when examining the mode-by-mode SNR of the 21 events we considered, 9 out of the 21 exhibited the above ordering. In Fig. 2, we can see the prevalence of this mode ordering within the region of parameter space we consider. For each inclination angle, different colors indicate a different ordering of the subdominant modes (ranked from “loudest” to “quietest”). In all the configurations shown here, the (2, 2) mode dominates, with the (3, 3) mode generally being the second loudest. The ordering of the (2, 1), (4, 4), and (3, 2) modes varies a bit more, with the (4, 4) usually being the quietest for a nearly face-on system, and the (3, 2) generally being the quietest otherwise.
@@
We use the power spectral density (PSD) required by the LISA Science Requirements Document (SciRDv1) [Babak:2021mhe]. We assume the frequency range of LISA to be $10^{-5}-0.5\,{\rm Hz}$ and incorporate the full response of LISA [Marsat:2018oam]. We also account for a background of white dwarf noise as in Sec. 9 of Ref. [Babak:2021mhe].
To sample the GW likelihood, we use the parallel-tempered Markov chain Monte Carlo (MCMC) ensemble sampler ptemcee [Vousden:2016eeu, Foreman-Mackey:2012any, Goodman:2010dyf], which is a fast sampler well-adapted to handling complex posteriors. To improve the robustness for degenerate posteriors, the ensemble sampler is enriched with jump proposals designed to explore the possible sky multimodality [Marsat:2020rtl]. For a typical PE run, we use 64 walkers and run ptemcee at 10 temperatures with a max temperature of 100. For each event, we sample in the combinations of mass and spin parameters $\mathcal{M}_{c},\,q,\,\chi_{+},$ and $\chi_{-}$] defined in Eq. (I). At first order in post-Newtonian (PN) theory, the GW frequency evolution is solely dependent on the chirp mass $\mathcal{M}_{c}$, so this is often the mass parameter that is most easily determined from the data. Similarly, $\chi_{+}$ (also commonly known as $\chi_{\rm{eff}}$) is generally much easier to determine from the data than the individual spins $\chi_{1}$ and $\chi_{2}$. Lastly, $q$ and $\chi_{-}$ are chosen to complete the set of mass and spin parameters, respectively.
To perform PE, we first calculate the Fisher covariance matrix and initialize chains from the results of this calculation. We set uniform priors on the chirp mass and distance in the range of 0.1 to 10 times their injected values. For the mass ratio, we use a uniform prior over the range [1, 10][1,\,10], and the spin parameters have a uniform prior over their allowed ranges of [1, 1][-1,\,1].
@@ II.2 Predicting biases in the linear signal approximation
Given the computational cost of full Bayesian PE analyses, it is desirable to have a much faster, cheaper method of predicting systematic biases due to, e.g., neglecting higher-order modes. One technique for accomplishing this is to estimate systematic errors in the linear signal approximation as done by Cutler & Vallisneri (henceforth, “CV”) in Ref. [Cutler:2007mi]. In Ref. [Antonelli:2021vwg], the errors calculated in this approximation were found to reliably predict the biases in parameter inference of ET and LISA sources due to overlapping signals, foregrounds created by unfitted sources, and incorrectly removed sources, among other things. The authors of Ref. [Pitte:2023ltw] examined the usefulness of this technique for the example binary they considered, finding that the errors were not properly reconstructed within the linear signal approximation. Here we expand the study of the usefulness of the CV estimates, covering a much larger region of parameter space than previously covered in Ref. [Pitte:2023ltw]. In Appendix B, for completeness, we review the CV formalism. In Sec. V below we compare our PE results with the errors predicted by the CV estimates.